Flight of the Hippogriff

Below is some information for Flight of the Hippogriff

Ticket price $25

Variable costs $6

Fixed costs 24,000

The theater contains 1,500 seats

The Board will like to make a profit of $40,000 per run

It is expected that you will have 3 shows and that each show will be about 70% full. For the full run of the show (all 3 shows) For each question below use the information given to answer the question.   Show all work

  1. Prepare a contribution margin income statement. What is the Net Operating Income for this show?

  2. What is the break-even point for this show in units (seats)? Compare this to the expected number of units sold. Explain how this impacts the theater’s profits.

  3. Will Flight of the Hippogriff meet the target profit set by the board?  

  4. What is the Margin of Safety?

  5. The performers in Flight of Hippogriff like to have a very full audience. They discuss dropping ticket prices to $20. They think the shows will be 80% full with a price decrease. Do they have a good idea? Explain why or why. Support your position with calculations.

  6. The board is thinking about spending an additional $1,500 on advertising. They estimate an increase of 1,000 seats sold. Do they have a good idea? Explain why or why. Support your position with calculations.
Answer #1

Flight of the Hippogriff


Contribution margin per ticket = $ 25 - $ 6 = $ 19

Expected ticket sales = 1,500 x 3 x 70 % = 3,150

6. Net operating income = 3,150 x $ 19 - $ 24,000 = $ 35,850

7. Break-even number of tickets = $ 24,000 / $ 19 = 1,263.16 tickets or 1,264 tickets

Expected tickets sold = 3,150

8. No.

9. Margin of safety = 3,150 - 1,264 = 1,886 tickets or $ 47,150

10. New contribution margin = $ 20 - $ 6 = $ 14

Expected ticket sales = 1,500 x 3 x 80 % = 3,600

Net operating income = 3,600 x $ 14 - $ 24,000 = $ 26,400

No, the idea of the performers is not good. Profit will decline by $ 9,450

11. Net operating income = ( 3,150 + 1,000 ) x $ 19 - $ 25,500 = $ 53,350.

Yes, the idea of the board is good, as the target profit will be achieved.

12.


Goblets of FireFlight of the Hippogriff
Sales$ 84,375$ 78,750
Variable costs50,62518,900
Contribution margin33,75059,850
Fixed costs2,00024,000
Net operating income31,75035,850
Degree of operating leverage ( Contribution margin / Net operating income )1.06301.669


answered by: hero